Assume that stock market returns do not resemble a single-index structure. An investment fund analyzes 100 stocks in order to construct a mean-variance efficient portfolio constrained by 100 investments. They will need to calculate ____________ covariances.

Multiple Choice

  • 4,950

  • 100

  • 45

  • 10,000
Answer #1

Answer:

Solution

Covariance= (n2-n)/2

=(1002-100)/2

=(10000-100)/2

=4950

Answer ; 4950